Circular No. 7151

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1999 H3 An apparently asteroidal object discovered by LINEAR, and noted on

The NEO Confirmation Page, has been reported as cometary by Kleť and Ondřejov observers. The following observations are available:

1999 UT α_{2000} δ_{2000} m_{1} Observer
Apr. 22.30901 19^{h}17^{m}19.^{s}76 +44^{o}09′48′′.1 16.8 Shelly
22.32527 19 17 18.66 +44 10 09.3 17.1  _{”}
22.33342 19 17 18.04 +44 10 20.0 17.0  _{”}
22.34245 19 17 17.36 +44 10 33.7 16.8  _{”}
23.87372 19 15 28.48 +44 44 32.2 15.5 Tichá
23.87542 19 15 28.31 +44 44 34.2  _{”}
23.87583 19 15 28.26 +44 44 33.8  Šarounová
23.87650 19 15 28.25 +44 44 35.6  Tichá
23.87781 19 15 28.13 +44 44 36.5  Šarounová

F. Shelly, M. Blythe, M. Bezpalko, and M. Elowitz (LINEAR). Measurers

J. Stuart, H. Viggh, and R. Sayer. 1.0-m f/2.15 reflector + CCD. J. Tichá and M. Tichý (Kleť). 0.57-m f/5.2 reflector + CCD. Coma diam-

eter ≥ 15′′, wide tail ≥ 13′′ in p.a. 166^{o}. L. Šarounová (Ondřejov). 0.65-m f/3.6 reflector + CCD. Coma diameter

> 20′′; poor observing conditions.

1999 CV_{3}

M. Kidger, Instituto de Astrofiscia de Canarias (IAC), communicates: “J. Altozano, J. Marti, I. Ugarte, and D. Hernandez, Astrophysics Department, La Laguna University, report that analysis of photometric observations of 1999 CV_{3} made with the 0.82-m IAC-80 Telescope at Teide Observatory on Mar. 5/6 and 6/7, together with observations by D. Buczinsky (Feb. 22, 0.33-m reflector, Condor Brow) and N. James (Feb. 13, 0.30-m reflector, Chelmsford), yields a best-fit lightcurve of period 5.78 ± 0.06 hr, with a double maximum of amplitudes 0.4 and 0.25 mag.”

COMET C/1999 H1 (LEE) Visual m_{1} estimates: Apr. 20.50 UT, 8.3 (S. T. Rae, Palmerston North,

N.Z., 10×50 binoculars); 21.58, 9.0 (M. Mattiazzo, Wallaroo, S. Australia, 0.20-m reflector); 22.54, 8.7 (A. Pearce, Nedlands, W. Australia, 0.20-m refl.); 22.96, 8.7 (J. G. de S. Aguiar, Campinas, Brazil, 0.23-m refl.).

1999 April 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green